A regression was run to determine if there is a relationship between hours of TV watched per day (x) and number of situps a person can do (y). The results of the regression were:

y=ax+b

a=-1.026 b=24.441 r2=0.958441 r=-0.979

Use this to predict the number of situps a person who watches 1 hours of TV can do (to one decimal place)


Please show work
Answer #1

y=ax +b

a=-1.026

b=24.441

y=ax+b

=-1.026x +24.441

We are given x=1

so

y=-1.026x +24.441

=-1.026(1) +24.441

=-1.026 +24.441

=23.415

=23.4